A mistake at the foundry has caused a 55.8 litre alloy of bronze to contain only 7% of tin instead of the necessary 13%.
How much tin needs to be added to the bronze to fix the mistake?

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

A mistake at the foundry has caused a 55.8 litre alloy of bronze to contain only 7% of tin instead of the necessary 13%.
How much tin needs to be added to the bronze to fix the mistake?
-------------------------
Add L liters of 100% tin.
55.8*7 + L*100 = (55.8 + L)*13
Solve for L

Answer by Theo(13342)

the amount of tin is currently .07 * 55.8 liters = 3.906 liters.

let x equal the amount of bronze that needs to be added.

your equation is:

3.906 + x = .13 * (55.8 + x).

simplify to get:

3.906 + x = 7.254 + .13 * x

subtract .13 * x from both sides of the equation and subtract 3.906 from both sides of the equation to get:

x - .13 * x = 7.254 - 3.906

combine like terms and simplify to get:

.87 * x = 3.348.

solve for x to get:

x = 3.348 / .87 = 3.848275862.

that's the amount of bronze that needs to be added to make the percentage of bronze equal to 13%.

you have 3.906 + 3.848275862 = 7.754275862 liters of bronze in a total of 55.8 + 3.48275862 = 59.64827586 liters.

the percentage of bronze is 7.754275862 / 59.64827586 = .13 * 100 = 13%.